Analysis of Receivables Method

At the end of the current year, Accounts Receivable has a balance of $735,000; Allowance for Doubtful Accounts has a debit balance of $6,500; and sales for the year total $3,310,000. Using the aging method, the balance of Allowance for Doubtful Accounts is estimated as $28,600.

a. Determine the amount of the adjusting entry for uncollectible accounts.
$

b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Bad Debt Expense.

Accounts Receivable $
Allowance for Doubtful Accounts $
Bad Debt Expense $

c. Determine the net realizable value of accounts receivable.
$
Answer #1

a) Adjusting entry amount = 28600+6500 = $35100

b) Account receivable = 735000

Allowance for doubtful accounts = 28600

bad debt expense = 35100

c) Net realizable value = 735000-28600 = 706400